# ANCHOLIK-NER: A Benchmark Dataset for Bangla Regional Named Entity Recognition

### v1

@@ Abstract
ANCHOLIK-NER is a linguistically diverse dataset for Named Entity Recognition (NER) in Bangla regional dialects, capturing variations across Sylhet, Chittagong, and Barishal. The dataset has around 10,443 sentences, 3,481 sentences per region. The data was collected from two publicly available datasets and through web scraping from various online newspapers, articles. To ensure high-quality annotations, the BIO tagging scheme was employed, and professional annotators with expertise in regional dialects carried out the labeling process. The dataset is structured into separate subsets for each region and is available both in CSV format. Each entry contains textual data along with identified named entities and their corresponding annotations. Named entities are categorized into ten distinct classes: Person, Location, Organization, Food, Animal, Colour, Role, Relation, Object, and Miscellaneous. This dataset serves as a valuable resource for developing and evaluating NER models for Bangla dialectal variations, contributing to regional language processing and low-resource NLP applications. It can be utilized to enhance NER systems in Bangla dialects, improve regional language understanding, and support applications in machine translation, information retrieval, and conversational AI.
@@
Data accessibility Repository Name: Mendeley Data (10.17632/gbkszkt8z3.1)
@@ 1 Value of the data
  • It includes text from multiple sources, with dialects such as Sylhet, Chittagong, and Barishal, enhancing its usability for dialect-aware NLP applications.
@@
  • Provided both in format, it includes textual data and named entity annotations, making it adaptable to different NER frameworks and machine learning models.
@@ 3 Data Description
We have developed the ANCHOLIK-NER, Bangla Regional Named Entity Recognition dataset, which focuses on the regional dialects of Sylhet, Chittagong, and Barishal. This dataset is specifically designed to enhance the recognition and classification of named entities within the context of these dialects, which are underrepresented in existing NER datasets. By capturing the linguistic nuances specific to these regions, the dataset provides a valuable resource for developing more accurate and regionally aware NER models for Bangla.
Our Bangla Regional Named Entity Recognition (NER) dataset comprises a total of 10,443 sentences, with 3,481 sentences collected for each of the three major regional dialects: Sylhet, Chittagong, and Barishal. Named entities in this dataset are systematically categorized into ten distinct types: Person, Location, Organization, Food, Animal, Colour, Role, Relationship, Object, and Miscellaneous. 71.81% of the dataset (sentences 1 to 2,500) of each region was sourced from two publicly available datasets (Vashantor[13] and ONUBAD[14]) and the rest 29.11% through web scraping from online newspapers and articles, covering structured and editorially curated texts. This incorporates informal, conversational, and spoken language characteristics unique to these regional dialects. This diverse data collection approach helps in capturing real-world linguistic variations and dialectal expressions, making the dataset more representative of actual language usage across different mediums. The number of instances of each type of named entities in the dataset for each three regions is given in Table 2.
@@
    A Python script was used to process the raw sentences, converting them into a structured tabular format. The script assigned a unique serial number to each sentence and extracted individual words into separate columns. This was done to convert the sentences into tokenized words for efficient annotation. The processed output was saved in both CSV format.
@@ 1 AaIech O
Table 2: Total instances of Named Entity Types in three Regions
Named Entity Type Barishal Sylhet Chittagonj Total Instances
Person (PER) 39 38 39 116
Location (LOC) 540 544 538 1622
Organization (ORG) 143 146 144 433
Food (FOOD) 310 308 308 926
Animal (ANI) 56 55 56 167
Colour (COL) 161 167 160 488
Role (ROLE) 117 110 112 339
Relation (REL) 664 660 662 1986
Object (OBJ) 348 345 345 1038
Miscellaneous (O) 17773 18590 18025 54388
@@ 4 Limitations
The dataset focuses on three regional dialects: Sylhet, Chittagong, and Barishal, potentially excluding linguistic variations present in other Bangla-speaking regions.

### v2

@@ Abstract
ANCHOLIK-NER is a linguistically diverse dataset for Named Entity Recognition (NER) in Bangla regional dialects, capturing variations across Sylhet, Chittagong, Barishal, Noakhali, and Mymensingh. The dataset has around 17,405 sentences, 3,481 sentences per region. The data was collected from two publicly available datasets and through web scraping from various online newspapers, articles. To ensure high-quality annotations, the BIO tagging scheme was employed, and professional annotators with expertise in regional dialects carried out the labeling process. The dataset is structured into separate subsets for each region and is available in CSV format. Each entry contains textual data along with identified named entities and their corresponding annotations. Named entities are categorized into ten distinct classes: Person, Location, Organization, Food, Animal, Colour, Role, Relation, Object, and Miscellaneous. This dataset serves as a valuable resource for developing and evaluating NER models for Bangla dialectal variations, contributing to regional language processing and low-resource NLP applications. It can be utilized to enhance NER systems in Bangla dialects, improve regional language understanding, and support applications in machine translation, information retrieval, and conversational AI.
@@
Data accessibility Repository Name: Mendeley Data (10.17632/gbkszkt8z3.2)
@@ 1 Value of the data
  • It includes text from multiple sources, with dialects such as Sylhet, Chittagong, Barishal, Noakhali, and Mymensingh, enhancing its usability for dialect-aware NLP applications.
@@
  • Provided in CSV format, it includes textual data and named entity annotations, making it adaptable to different NER frameworks and machine learning models.
@@ 3 Data Description
We have developed the ANCHOLIK-NER, Bangla Regional Named Entity Recognition dataset, which focuses on the regional dialects of Sylhet, Chittagong, Barishal, Noakhali, and Mymensingh. This dataset is specifically designed to enhance the recognition and classification of named entities within the context of these dialects, which are underrepresented in existing NER datasets. By capturing the linguistic nuances specific to these regions, the dataset provides a valuable resource for developing more accurate and regionally aware NER models for Bangla.
Our Bangla Regional Named Entity Recognition (NER) dataset comprises a total of 17,405 sentences, with 3,481 sentences collected for each of the three major regional dialects: Sylhet, Chittagong, Barishal, Noakhali, and Mymensingh. Named entities in this dataset are systematically categorized into ten distinct types: Person, Location, Organization, Food, Animal, Colour, Role, Relationship, Object, and Miscellaneous. 86.18% of the dataset (sentences 1 to 3000) of each region was sourced from two publicly available datasets (Vashantor[13] and ONUBAD[14]) and the rest 13.12% through web scraping from online newspapers and articles, covering structured and editorially curated texts. This incorporates informal, conversational, and spoken language characteristics unique to these regional dialects. This diverse data collection approach helps in capturing real-world linguistic variations and dialectal expressions, making the dataset more representative of actual language usage across different mediums. The number of instances of each type of named entities in the dataset for each three regions is given in Table 2.
@@
    A Python script was used to process the raw sentences, converting them into a structured tabular format. The script assigned a unique serial number to each sentence and extracted individual words into separate columns. This was done to convert the sentences into tokenized words for efficient annotation. The processed output was saved in CSV format.
@@ 1 AaIech O
Table 2: Total instances of Named Entity Types in five Regions
Named Entity Type Barishal Sylhet Chittagong Noakhali Mymensingh Total Instances
Person (PER) 39 38 39 39 39 194
Location (LOC) 369 371 377 361 362 1840
Organization (ORG) 139 141 139 141 140 700
Food (FOOD) 310 308 308 303 312 1541
Animal (ANI) 57 56 57 57 57 284
Colour (COL) 162 167 160 164 163 816
Role (ROLE) 114 107 109 111 113 554
Relation (REL) 681 677 676 676 676 3386
Object (OBJ) 352 348 348 350 349 1747
Miscellaneous (O) 17928 18750 18177 17957 17943 90,755
@@ 4 Limitations
The dataset focuses on only five regional dialects: Sylhet, Chittagong, Barishal, Noakhali, and Mymensingh, potentially excluding linguistic variations present in other Bangla-speaking regions.